### v2

# Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects

Abstract
Novel metaphor comprehension involves complex semantic processes and linguistic creativity, making it an interesting task for studying language models (LMs). This study investigates whether surprisal, a probabilistic measure of predictability in LMs, correlates with different metaphor novelty datasets. We analyse surprisal from 16 LM variants on corpus-based and synthetic metaphor novelty datasets. We explore a cloze-style surprisal method that conditions on full-sentence context. Results show that LMs yield significant moderate correlations with scores/labels of metaphor novelty. We further identify divergent scaling patterns: on corpus-based data, correlation strength decreases with model size (inverse scaling effect), whereas on synthetic data it increases (Quality–Power Hypothesis). We conclude that while surprisal can partially account for annotations of metaphor novelty, it remains a limited metric of linguistic creativity. [^1] Code and data are publicly available: https://github.com/OmarMomen14/surprisal-metaphor-novelty.
Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects
Omar Momen,   Emilie Sitter, Berenike Herrmann and Sina Zarrieß CRC 1646 – Linguistic Creativity in Communication Faculty of Linguistics and Literary Studies Bielefeld University, Germany {omar.hassan,emilie.sitter,berenike.herrmann,sina.zarriess}@uni-bielefeld.de
1 Introduction
Recent advances in neural language modelling have led to a renewed interest in studying linguistic creativity, an aspect of language that was notoriously challenging for traditional NLP systems. A well-known instance of creative language is metaphor, which arises through mapping of a source domain–a lexical unit’s literal meaning–onto a target domain–its figurative meaning-Lakoff and Johnson (1980). Yet, not all metaphors are equally novel or creative as many of such mappings are highly conventionalized, such as in the famous example “to attack an argument,” where the mapping from the source domain “WAR” onto the target domain “ARGUMENT” corresponds to a conventional sense of the verb “attack” which can be found in dictionaries. Other metaphors, as in “The arrested water” (Figure 1) establish a more unconventional or novel mapping that is considered creative.
Figure 1: A sentence from VUA-ratings with metaphor-novelty ratings (n) and surprisal measures (s) from GPT2-base. The top part shows the direct-surprisal measure, while the bottom part shows the cloze-surprisal measure.
The distinction between conventional and novel metaphors is well established in theory Bowdle and Gentner (2005); Gibbs (2005) and studied in experimental work on cognitive processing Arzouan et al. (2007); Lai et al. (2009). Novel metaphors require greater interpretative effort compared to conventional metaphors, as the unfamiliar mapping requires speakers to construct new connections between domains Philip (2016).
However, separating novel metaphors from conventional ones remains a challenging task for human-based annotations Do Dinh et al. (2018); Reimann and Scheffler (2024). While computational methods on metaphor detection and understanding do not even take this distinction into account in most cases (Maudslay and Teufel, 2022), and some existing studies highlight that novel metaphors are more difficult to detect than conventional ones Neidlein et al. (2020); Tong et al. (2024). In this work, we propose to fill this gap and study questions on LMs processing of novel and conventional metaphors, and investigate LM-based metrics correlations to different setups of metaphor novelty annotations.
We draw on a line of research that has studied the ability of LMs to account for effects of difficulty in human sentence processing and that goes back to surprisal theory (Hale, 2001; Levy, 2008). Surprisal is computed with LMs as the negative log probability of a word in context and has been found to provide a robust predictor for human processing difficulty (e.g., of reading times) (Goodkind and Bicknell, 2018; Shain et al., 2024).
However, recent work draws a mixed picture in terms of which LMs can provide the most robust and cognitively plausible predictors. Oh and Schuler (2023b) observe an inverse scaling effect when testing surprisal estimates from GPT-2 models of different sizes, showing that the surprisal computed with smaller model sizes achieved a better fit with human reading times than larger model sizes. Wilcox et al. (2023), on the other hand, train LMs of small and medium size on a range of languages and find that LM quality generally correlates with its psychometric predictive power.
In this paper, we investigate surprisal on metaphor words computed with LMs as a metric of metaphor novelty. We find significant moderate correlations between LMs’ surprisal and different metaphor novelty annotations. We perform our experiments on four different metaphor novelty datasets coming from corpus-based and synthetic setups, using LMs of different sizes and architectures. Most interestingly, we observe effects supporting the "inverse scale" pattern Oh and Schuler (2023b) on the two corpus-based novelty datasets, and contrary patterns supporting the counter-argument, "Quality-Power" hypothesis Wilcox et al. (2023) on the two synthetic datasets. We also investigate the effect of instruction-tuning on surprisal correlation to metaphor novelty. Moreover, we conduct a deeper analysis of the genre splits, revealing that genre, metaphor density and LM perplexity are potential factors underlying the quality of surprisal as a predictor of metaphor novelty. Finally, we introduce a new method of computing surprisal, cloze-surprisal, to include the right context of the word in its conditional probability. We find that this method can boost the correlation of surprisal with metaphor novelty annotations by a few points. In general, our study establishes a promising direction for studying linguistic creativity with LMs and calls for novel measures and datasets that provide systematic annotations of metaphor novelty across genres.
2 Related Work
2.1 Metaphor Annotations
The most common annotation scheme in corpus-based metaphor studies is the Metaphor Identification Procedure Vrije Universiteit (MIPVU) The Pragglejaz Group (2007); Steen et al. (2010), designed as a reliable step-by-step framework for identifying metaphorically used words. This method was used to construct the VU Amsterdam Metaphor Corpus (VUAMC) Steen et al. (2010), a large-scale, genre-balanced corpus containing 186,673 words sampled from fiction, news, academic writing, and conversations splits in the BNC Baby edition. The VUAMC has become a benchmark for metaphor research, inspiring developments in further English corpora Beigman Klebanov et al. (2018); Mohammad et al. (2016) and more recent multilingual efforts Sanchez-Bayona and Agerri (2022); Egg and Kordoni (2022).
Several studies have extended annotation beyond binary literal/metaphoric labels. Mohler et al. (2016) introduced a four-point metaphoricity scale, considering factors such as vividness and familiarity. Reijnierse et al. (2019) proposed the concept of deliberateness, distinguishing metaphors intended to be recognized as metaphors; novel metaphors are typically deliberate under this framework. Direct novelty annotations were first introduced by Parde and Nielsen (2018), who asked annotators to rate metaphorical word pairs from the VUAMC on a 0–3 novelty scale. Do Dinh et al. (2018) proposed a more comprehensive approach considering all metaphor words in VUAMC, and by aggregating ranked novelty judgments into continuous scores. More recently, Reimann and Scheffler (2024) proposed a dictionary-based method that labels a metaphor as novel when its contextual meaning is absent from dictionary entries.
Psycholinguistic studies also produce datasets of metaphors, but these are smaller in size and feature controlled synthetic sentences (fixed words, structures, sentence length, etc.), cf. Cardillo et al. (2010, 2012); Roncero and de Almeida (2014). Some of these also created datasets of metaphors that are classified as conventional or novel metaphors Lai et al. (2009); Ahrens et al. (2024).
2.2 Predictive Powers of Surprisal
Surprisal estimates from LMs have been shown to capture human processing difficulty across multiple behavioural and neural measures. In self-paced reading and eye-tracking studies, token-level surprisals from causal models significantly predict reading times beyond lexical and syntactic factors Goodkind and Bicknell (2018); Wilcox et al. (2020); Oh and Schuler (2023a). Similar effects have been reported for acceptability judgments, where LM surprisals align with human sensitivity to gradient grammaticality Lau et al. (2017); Meister et al. (2021); Hu et al. (2020); Tjuatja et al. (2025).
2.3 LM-Based Metrics for Metaphor Novelty
Recent work has explored diverse LM-based approaches—prompting, surprisal, embedding similarity and attention patterns—for the analysis of metaphorical language. Aghazadeh et al. (2022) demonstrate that metaphor-relevant information is encoded in mid-layer embeddings of multilingual pre-trained LMs. Ichien et al. (2024) show that GPT-4 generates interpretations of novel literary metaphors that are favoured by human judges over interpretations by college students, suggesting sensitivity to metaphorical meaning beyond lexical overlap. Pedinotti et al. (2021) tested BERT’s masked token probabilities across conventional, novel, and nonsensical metaphors, showing that novel metaphors tend to receive lower probabilities, but distinctions between novelty and deviancy remained unclear. We extend this study towards more recent LMs and broader comparisons between models, datasets, and genres. Djokic et al. (2021) trained a BERT-based classifier to predict novelty scores jointly with the task of metaphor detection.
3 Metaphor Novelty Datasets
This Section describes the four datasets we use in our experiments and explains how metaphor novelty can be annotated in different approaches. Datasets statistics are reported in Tables 6 and 7.
3.1 Corpus‐based Datasets
Following the MIPVU procedure, all words in the VUAMC are annotated as being a metaphor-related word (MRW) or not. Out of 186,673 words in VUAMC, 24,762 words (15,155 content words) are labelled as MRWs. While the original VUAMC does not include novelty annotations, we utilise two annotation studies that offer different annotation protocols for metaphor novelty on the same corpus.
VUA-ratings
Do Dinh et al. (2018) collected crowd-sourced ratings of metaphor novelty for VUAMC. In their set-up, annotators were presented four sentences containing an MRW (content words only) and asked to select the best (most novel) and worst (most conventional). Each MRW appeared in six different best-worst scaling comparisons. These annotations were then converted into continuous Best-Worst Scaling scores Kiritchenko and Mohammad (2017) in the range of (-1, +1), with -1 being the most conventional and +1 being the most novel. Additionally, they also convert these scores to binary labels using a threshold of 0.5. This results in labelling 353 metaphors as novel out of 15,155 content metaphoric words in VUA.
VUA-dictionary
Reimann and Scheffler (2024) proposed a dictionary-based method that labels a metaphor as novel when its contextual meaning is absent from dictionary entries. They applied this method to VUAMC. In particular, they re-annotated the potentially novel metaphors in VUA according to VUA-ratings and Reijnierse et al. (2019)’s metaphor deliberateness annotations (1,160 potentially novel metaphors in total). Their procedure resulted in labelling only 409 content [^2] We exclude non-content words from this dataset to allow comparison with VUA-ratings, which originally excluded non-content words. metaphoric words as novel out of the 1,160 potentially novel metaphors. We assume the remaining metaphors in VUA to be conventional in our study.
3.2 Synthetic Datasets
Another class of metaphor novelty datasets is synthetic datasets, which are mainly characterised by being generated from a fixed source. They usually have comparable sentences in terms of the target metaphoric words. We consider two cases of such datasets, a dataset from a psycholinguistic study concerned with novel metaphors. And a toy dataset that we generated from GPT-4o.
Lai2009
Lai et al. (2009) investigated how our brains handle conventional and novel metaphors differently. To this end, they constructed a set of controlled experimental items featuring metaphors with two degrees of novelty. Two linguists selected 104 words, and constructed 4 sentences for each word, according to the Conceptual Metaphor Theory (CMT) Lakoff and Johnson (1980). For each word, the items include one (i) literal use, (i) conventional metaphor, (iii) novel metaphor and (iv) anomalous use, with the target word as the last word in each sentence. Familiarity and interpretability tests showed a significant difference between the conventional and novel metaphoric senses. For our experiments, we select only the conventional and novel metaphor senses for each word, resulting in 208 sentences (104 conventional and 104 novel metaphors).
GPT-4o-metaphors
To test our experiments in a setting that is more controlled (in contrast to VUA) but features more varied sentence lengths, structures and degrees of novelty (in contrast to Lai2009), we construct a synthetic dataset by prompting GPT-4o to generate sentences that include conventional and novel metaphoric senses. We do not assume that LLMs are able to generate ideal novel metaphors, or that this dataset is a benchmark of any kind; we are only interested in exploring a new quick setting of potential metaphor novelty annotations. We prompt GPT-4o to generate 5 verbs and 5 nouns that can be used in a metaphoric sense. For each word, we prompt the model again to generate 10 different sentences using the target word in a conventional metaphor sense, and 10 different sentences in a novel metaphor sense. This results in a dataset of 200 sentences, 100 contain conventional metaphors, and 100 contain novel metaphors.
4 Methods & Experiments
We describe our experiments designed to investigate whether and to what extent surprisal scores computed from LMs correlate with different annotations of metaphor novelty.
4.1 Surprisal of a target word
In information theory Shannon (1948), the information content of an event $x$ with probability $p(x)$ is defined as: $I(x)=-\log p(x)$.
In the context of LMs trained to predict the next token in a sequence, this quantity is called surprisal and computed for a $w_{i}$ in a sequence such that: $\mathrm{Surprisal}(w_{i})=-\log p(w_{i}\mid w_{<i})$ [^3] We use $\log$ of base $e$ for all surprisal/perplexity computations in our study..
In our experiments, we measure word-level surprisal of metaphoric words in their sentence-level context. We feed every sentence to an LM in an independent teacher-forced forward pass, and record surprisal of the target (metaphoric) word(s). We denote this quantity as direct-surprisal to distinguish from cloze-surprisal.
Naturally, the method of computing word-level surprisal from LMs is a rough approximation process, and design choices of this method can bias the results Pimentel and Meister (2024); Lesci et al. (2025). We identify a couple of choices in the details of this method that can affect  [^4] We find differences in the correlation values when these choices are altered. the final results of our study. Mainly, subword tokenisation is the reason for such methodological confounds.
Firstly, the goal is to get the probability of a target word (e.g. “arrested” in Figure 1); however, most LMs produce probabilities per tokens (subwords) (e.g. [‘‘Ġarrest’’, ‘‘ed’’]) rather than words. And the method of aligning these tokens from the list of sentence tokens to specific words becomes a methodological choice. We choose to align the tokens to words by precomputing the offsets of the exact [^5] We make sure no punctuation or any other characters outside of the word boundary are included inside the offsets. target words in their corresponding sentences, and then searching for the minimal span of tokens in the LM’s tokenisation of the input sentence covering the target word’s offsets. This usually results in a proper alignment (e.g. ‘‘Ġarrested’’ for “arrested”), with very rare cases when the last token is attached to a punctuation (e.g. [‘‘Ġindivid’’, ‘‘ual’’, ‘‘ism,’’] for “individualism”). We also apply the corrections made by Pimentel and Meister (2024); Oh and Schuler (2024) that address the problem of leading whitespaces in tokenisations of most causal LMs.
Secondly, surprisal is computed using the conditional probability of the current word being the target word given the preceding words in the sentence. This makes computing surprisal of the first word in a sentence an issue [^6] In some cases in the VUA datasets, the target word is the first word in the sentence.. To allow computing surprisal of the first word in a sentence, we prepend the input sentence with a “begining of sequence” special token.
Cloze-surprisal
We try to address this inherent weakness of word-level surprisal computed from causal models, which does not take the right context of the metaphoric word into account. This is an issue when sentences do not end in the metaphorical word, but feature various positions of the metaphorical word as in naturally occurring corpus data (see Figure 4), and also in some cases of the GPT-4o-metaphors dataset. To address this, we experiment with a new strategy for computing word-level surprisal autoregressively, while introducing the full sentence context: for each metaphor word, we input a prompt (as illustrated in Figure 1) that indicates a cloze test and show the whole sentence to the model except the metaphor word (replaced by a blank space), and then spell out the whole sentence again as if the model is solving the cloze test. We measure the surprisal of the target word at the second occurrence of the sentence, ensuring that the left and right context are included in the conditional probability of the metaphor word.
4.2 Evaluating the correlation
To determine whether LMs surprisal correlates with different metaphor novelty annotations, we measure multiple correlation metrics between surprisal and metaphor novelty scores/labels. For continuous novelty scores (as in VUA-ratings), we compute Pearson’s $r$ Pearson (1895) and Spearman’s $\rho$ Spearman (1904) correlations. For binary labels (conventional vs. novel), we compute the Rank-biserial $r_{b}$ Glass (1966) correlation. Rank-biserial estimates the probability that a random observation from the set of novel metaphors has a larger surprisal than one from the set of conventional metaphors, minus the reverse probability. We also estimate the potential of surprisal as a discriminator for binary novelty labels using the Area Under the ROC Curve (AUC) Fawcett (2006). We also compute the significance of all these estimates. All our metrics and tests do not assume normality, except for Pearson’s, for which we also provide a non-parametric alternative (Spearman).
4.3 Settings
Surprisal measures are derived from the learnt $p(x)$ of pre-trained LMs and, hence, depend substantially on the model architecture and the training data. Also, human annotations of metaphor novelty depend on the annotation process. Additionally, novelty norms may vary by genre: a metaphor considered novel in academic texts might appear conventional in fiction. We thus investigate how all these factors affect the correlation by experimenting with multiple settings.
Models
We examine three families of decoder-only causal LMs (GPT-2, Llama 3 and Qwen2.5) Radford et al. (2019); Grattafiori et al. (2024); Yang et al. (2024). For each model family, we select 3-4 different sizes, to represent the effect of model size on the correlation. To investigate the effect of instruction-tuning on the correlation, we include the instruction-tuned variants of Llama 3 and Qwen2.5.
Datasets
As explained in Section 3, we perform experiments on four datasets (Tables 6 and 7). As VUA-ratings has continuous novelty scores, we evaluate the correlation of its continuous scores to surprisal. In addition, we convert these scores to binary labels using a threshold of $0.5$ and evaluate their correlation to surprisal, allowing us to compare the results of VUA-ratings to VUA-dictionary, Lai2009 and GPT-4o-metaphors, which only have binary novelty annotations.
Genre variables
VUA provides genre splits (see Section 3.1), allowing us to analyse the correlation between surprisal and novelty separately for each genre.
Perplexity
Scales of surprisal can differ from one genre to another. Perplexity of a model on a certain dataset is the exponential of the average token-level surprisal of all tokens in the dataset. In that sense, perplexity indicates how much a certain model is “surprised” on average when predicting a certain dataset. We measure models’ perplexity on the genre splits of VUA by feeding sentences one by one to the model, measuring token-level surprisals accordingly and averaging and exponentiating them to obtain perplexity. This yields higher values of perplexity than common values reported in literature, as we use a shorter context (single sentences).
4.4 Summary
Our experiments rely on a large collection of surprisal measures for each metaphor word in the four datasets: 32 different surprisal values (including direct and cloze) for 15,155 metaphoric words of the VUA dataset and for 208 and 200 metaphoric words of Lai2009 and GPT-4o-metaphors, respectively. In Figure 3, we plot the distributions of surprisal values by model, in addition to metaphor novelty scores/labels from VUA. From these plots and further normal distribution tests, we find that novelty scores and surprisal values are not strictly normally distributed.
5 Results
We report the correlation measures for the corpus-based datasets in Table 1, and for synthetic datasets in Table 2. We also report the gains (in terms of rank-biserial) of the instruction-tuned variants over their base models in Table 3. And the gains of the cloze-surprisal method over the direct-surprisal method in Table 4.
5.1 Overall Correlation Results
Generally, we find LMs’ direct surprisal values correlate positively with metaphor novelty annotations across the four datasets (Tables 1 and 2). All correlation estimates are statistically significant for both the corpus-based and synthetic datasets  [^7] The relatively large corpus-based datasets (15,155 datapoints) can bias the significance values; however, we get similar significance values for the relatively small synthetic datasets (100 datapoints).. Overall for direct surprisal, the largest Pearson correlation $r=.419$, largest Spearman correlation $\rho=.417$, largest rank-biserial correlation $r_{b}=.638$ and largest $AUC=.819$ come from GPT2-base on VUA-ratings.
We also find that the correlation strengths’ scales differ across the different annotation datasets. By comparing the rank-biserial estimate $r_{b}$ across the four datasets, we find its ranges as follows: (.47-.64) for VUA-ratings, (.43-.58) for VUA-dictionary, (.28-.50) for Lai2009 and (.38-.63) for GPT-4o-metaphors. These results show that surprisal correlates relatively more with the corpus-based datasets than the synthetic datasets. Also, surprisal correlates relatively more with human ratings on VUA than the dictionary-based annotation approach. While the most controlled dataset (Lai2009) gets the weakest correlations with surprisal.
Model VUA-ratings VUA-dict.
$r$ $\rho$ $r_{b}$ aucauc $r_{b}$ aucauc
GPT2-base .419 .417 .638 .819 .581 .791
GPT2-med .389 .383 .600 .800 .557 .778
GPT2-large .381 .373 .585 .793 .539 .769
GPT2-xl .373 .362 .566 .783 .539 .769
Llama-1B .345 .329 .532 .766 .480 .740
Llama-3B .328 .308 .502 .751 .446 .723
Llama-8B .314 .293 .488 .744 .431 .716
Qwen-0.5B .384 .377 .598 .799 .543 .771
Qwen-7B .334 .314 .502 .751 .456 .728
Qwen-14B .316 .295 .470 .735 .430 .715
Table 1: Correlation estimates between surprisal and novelty scores/labels in the corpus-based datasets. All results are statistically significant at the 0.001 level.
Model Lai2009 GPT-4o-met.
$r_{b}$ aucauc $r_{b}$ aucauc
GPT2-base .276 .638 .511 .756
GPT2-med .362 .681 .586 .793
GPT2-large .397 .699 .629 .814
GPT2-xl .414 .707 .587 .794
Llama-1B .450 .725 .508 .754
Llama-3B .451 .725 .511 .755
Llama-8B .483 .742 .557 .778
Qwen-0.5B .374 .687 .382 .691
Qwen-7B .494 .747 .469 .734
Qwen-14B .504 .752 .536 .768
Table 2: Correlation estimates between surprisal and novelty labels in the synthetic datasets. All results are statistically significant at the 0.001 level.
Model VUA-r VUA-d Lai GPT-4o
Llama-1B-It. +4.0+4.0 +1.5+1.5 +0.4+0.4 5.1-5.1
Llama-3B-It. +4.0+4.0 +2.4+2.4 5.8-5.8 5.6-5.6
Llama-8B-It. +1.6+1.6 +0.6+0.6 2.3-2.3 5.1-5.1
Qwen-0.5B-It. 1.6-1.6 0.8-0.8 +0.1+0.1 +2.2+2.2
Qwen-7B-It. 1.7-1.7 1.9-1.9 8.9-8.9 13.7-13.7
Qwen-14B-It. 3.9-3.9 3.1-3.1 3.7-3.7 14.0-14.0
Table 3: Instruction-tuning % gains (over corresponding base variants) in Rank-biserial’s correlation estimates between surprisal and novelty scores/labels in the four datasets: VUA-ratings, VUA-dictionary, Lai2009 and GPT-4o-metaphors.
5.2 Model Size Effects
The effect of model size is consistent (but diverging in direction) across the two dataset types. We plot the rank-biserial correlations of the ten model variants for the four datasets in Figure 2. The two corpus-based datasets (blue) show a monotonically decreasing correlation as model size increases per model family. On the other hand, the two synthetic datasets (red) exhibit an increasing correlation behaviour as model size increases, with a single minor exception at the GPT2 model family on the GPT-4o-metaphors dataset.
Figure 2: Effect of model size on correlation between surprisal and metaphor novelty annotations from the four datasets. Corpus-based datasets (blue) show a negative scale effect, while synthetic datasets (red) show a positive scale effect.
5.3 Instruction-tuning Effects
We report the percentage gains in rank-biserial correlation for the instruction-tuned variants over their base variants across the four datasets in Table 3. To avoid confusion, these results examine the effect of extracting surprisals from an instruction-tuned variant over a base variant with the same experimental setting. We do not add a prompt/instruction to the instruction-tuned variants inputs. The results show that instruction-tuning does not always improve the correlation between surprisal and novelty annotations. Only Llama instruction-tuned variants could improve the correlations on the VUA datasets; otherwise, instruction-tuning deteriorates the correlations.
5.4 Cloze-surprisal
In Table 4, we report the percentage gains in rank-biserial correlations for the cloze-surprisal values over the direct-surprisal values from the same model variants across the four datasets. We find cloze-surprisal to be boosting the correlations with extra points in many cases; however, it also deteriorates the correlations heavily in other cases. In Figure 3, we can observe the distribution of cloze-surprisal of the metaphoric words shifted away from zero and approaching a normal distribution. Also in Figure 4, we observe the same effect. However, against our intuition, the effect of cloze-surprisal is not more impactful in cases where the metaphor is located earlier in the sentence than later. The effect of cloze-surprisal, although positive, is not related to the position of the metaphoric word in the sentence.
For the GPT2 model family, cloze-surprisal consistently improves the corpus-based datasets’ correlations, achieving the strongest correlations in this study (out of the 32 recorded surprisals on each dataset) coming from GPT2-base on VUA-ratings with values of $r=.499$, $\rho=.499$, $r_{b}=.687$ and $auc=.843$. However, it deteriorates the correlations at large scales on the synthetic datasets.
Llama and Qwen model families also show positive gains in most of the cases. Most importantly, unlike the GPT2 model family, it boosts correlations on the synthetic datasets except for a few cases. Cloze-surprisal achieves the overall strongest correlation on the Lai2009 dataset using the Qwen-14B model, with $r_{b}=.539$ and $auc=.779$. Also, it achieves the overall strongest correlation on the GPT-4o-metaphors dataset using the Llama-8B model, with $r_{b}=.684$ and $auc=.842$.
Model VUA-r VUA-d Lai GPT-4o
GPT2-base +4.9+4.9 +4.0+4.0 7.7-7.7 16.7-16.7
GPT2-med +6.8+6.8 +4.5+4.5 11.8-11.8 26.9-26.9
GPT2-large +6.2+6.2 +8.2+8.2 14.2-14.2 25.1-25.1
GPT2-xl +6.3+6.3 +2.5+2.5 16.0-16.0 28.1-28.1
Llama-1B +7.3+7.3 +8.0+8.0 13.3-13.3 15.7-15.7
Llama-1B-It. 3.8-3.8 1.2-1.2 2.4-2.4 10.3-10.3
Llama-3B 0.5-0.5 +5.6+5.6 +0.3+0.3 +0.7+0.7
Llama-3B-It. 11.1-11.1 2.3-2.3 +4.7+4.7 +6.5+6.5
Llama-8B +1.9+1.9 +2.7+2.7 +2.5+2.5 +12.7+12.7
Llama-8B-It. 7.8-7.8 0.2-0.2 +6.0+6.0 +10.9+10.9
Qwen-0.5B 2.1-2.1 +2.1+2.1 2.5-2.5 +6.2+6.2
Qwen-0.5B-It. 4.5-4.5 1.6-1.6 3.9-3.9 +7.0+7.0
Qwen-7B +3.7+3.7 +5.6+5.6 +2.4+2.4 +12.8+12.8
Qwen-7B-It. 6.9-6.9 1.6-1.6 +9.1+9.1 +8.0+8.0
Qwen-14B +5.8+5.8 +7.8+7.8 +3.5+3.5 +6.9+6.9
Qwen-14B-It. 4.4-4.4 +0.7+0.7 +4.6+4.6 +12.8+12.8
Table 4: Cloze-surprisal % gains (over direct-surprisal from the same model) in Rank-biserial’s correlation estimates between surprisal and novelty scores/labels in the four datasets: VUA-ratings, VUA-dictionary, Lai2009 and GPT-4o-metaphors.
Genre VUA-ratings VUA-dictionary
ppl. Nov. % $r_{b}$ Nov. % $r_{b}$
Fiction 108 2.97 .693 3.56 .478
News 89 2.80 .653 3.88 .413
Academic 73 1.85 .588 1.85 .413
Conversation 134 1.41 .482 0.62 .800
All 96 2.33 .638 2.70 .581
Table 5: GPT2-base perplexity, metaphor density, and GPT2-base surprisal correlation with metaphor novelty of VUA-ratings and VUA-dictionary per genre.
5.5 Genre Splits
In Table 5, we report the rank-biserial correlations between GPT2-base surprisals and binary novelty labels of VUA-ratings and VUA-dictionary on each genre split separately. We also report the share of novel metaphors within each genre to illustrate the variance in metaphor density across genres. Additionally, we report the perplexity of GPT2-base on each genre split separately.
For VUA-ratings, we find a positive relation between genres’ metaphor density and the correlation of surprisal with novelty. We also observe a positive relation between perplexity and surprisal-novelty correlations that is only violated on the conversation split. On the other hand, for VUA-dictionary, the relation between metaphor density and surprisal-novelty correlation is not preserved. While the perplexity positive relation with surprisal-novelty correlation is more concrete. Interestingly, in contrast to VUA-ratings, the dictionary-based annotations of the conversation split correlate strongly with surprisal values despite its tiny number of annotated novel metaphors (11 out of 1774 metaphors) and high perplexity (134).
6 Discussion
Is surprisal a good metric for metaphor novelty?
Across multiple novelty annotation setups—continuous human ratings, dictionary-based binary labels, linguist-designed contrasts, and LLM-generated conventional vs. novel senses—we find consistently moderate associations between LM surprisal and metaphor novelty (best: $r=.49,\rho=.50,r_{b}=.69,\mathrm{AUC}=.84$). While these results are not directly comparable to prior work due to differences in datasets and task formulations, their magnitudes are broadly in line with reported surprisal–behavior associations in acceptability judgments Tjuatja et al. (2025) and reading times Oh and Schuler (2023a). The closest point of reference in the metaphor novelty literature is Parde and Nielsen (2018), who report $r=.44$ when predicting novelty scores from a wide range of linguistic features. They additionally report moderate inter-annotator agreement, underscoring that novelty itself is challenging to measure reliably by humans.
At the same time, surprisal has clear theoretical limits as a standalone predictor: interpreting a metaphor involves semantic integration and cross-domain mapping, not only predictability. We therefore expect surprisal to be most informative when combined with measures that more directly target novel interpretations and domain mappings, especially for highly creative (novel) metaphors.
Model Sizes & Dataset Types:
Our results introduce a great opportunity to further understand the underlying factors of opposing negative and positive effects of model sizes on correlations with psycholinguistic and cognitive features. The negative and positive effects are present together in our study, and clearly contrasted by the type of metaphor novelty dataset under experiment.
The negative effect of model size observed in corpus-based datasets mirrors the inverse scaling effect reported in reading time studies Oh and Schuler (2023b); Wilcox et al. (2025). Recently, Oh et al. (2024) argued that this effect is largely driven by word frequency, with larger models assigning increasingly non-human-like expectations to rare words. Since corpus-based novelty scores are often correlated with word frequency Do Dinh et al. (2018); Reimann and Scheffler (2024), frequency can be an underlying factor for this negative scaling effect.
By contrast, the synthetic datasets control for lexical identity and frequency: conventional and novel senses are elicited for the same set of words, and stimuli are constructed to cleanly separate the two senses. Under these controls, scaling improves alignment with novelty labels, agreeing with reports of positive scaling effects in other behavioural settings such as acceptability judgments and other reading time studies Tjuatja et al. (2025); Wilcox et al. (2023).
Overall, we suggest that these diverging scaling effects are due to the nature of the dataset types. While corpus-based datasets reflect metaphor novelty mainly through lexical properties such as word frequency, synthetic datasets more directly isolate the conventional–novel distinction by controlling for lexical properties confounds.
Cloze-surprisal and Instruction-tuning:
Our cloze-surprisal approach improves correlation across many model variants. In Figures 3 and 4, cloze-surprisal is found to be shifting the surprisal values away from near-zero values and pushing towards a normal distribution.
In examples from Table 8, cloze-surprisal often raises the surprisal of the metaphorical word, consistent with the intuition that the full context is needed to process the metaphor and realise its true predictability. Moreover, in cases where the metaphorical word begins a sentence (e.g. example 14), direct surprisal is naturally high—regardless of novelty—whereas cloze-surprisal successfully moderates such inflated values.
Instruction-tuning—despite its goal of aligning model outputs with human intent—does not enhance the human-likeness of surprisal. In fact, similar to prior findings Kuribayashi et al. (2024), instruction-tuned models tend to reduce alignment between predicted probabilities and human annotation scores.
Genre Effects:
Our results show a significant effect of sentences’ genre on surprisal correlation with metaphor novelty. We suspect the difference in genre’s novel metaphor density can be an underlying factor. Also, our observations on perplexity relations to the correlations trigger the possibility that the amount of pretraining data contributing to each genre can significantly affect LM-based methods of detecting novel metaphors.
7 Conclusion
We have studied the distinction between conventional and novel metaphors and systematically investigated surprisal computed with LMs as a metric for metaphor novelty. In general, our experiments show some potential for surprisal in predicting aspects of linguistic creativity, but also call for novel measures and datasets that provide systematic annotations of metaphor novelty across genres and across corpus-based and experimental settings.
Limitations
We acknowledge several limitations in this work. First, the scarcity of high-quality metaphor novelty annotations in existing literature constrains both coverage and generalizability. Second, we rely on pretrained language models whose training data and processes are not fully disclosed.
Finally, although we carefully analyse the effects of model architecture, size, and domain (genre), future work could adopt mixed-effects models to test the interaction of these variables.
Acknowledgments
This research has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – CRC-1646, project number 512393437, project A05.
We acknowledge the reviewers for their valuable comments and feedback. And we also thank Vicky Tzuyin Lai and Tiago Pimentel for their responsiveness to our inquiries.
Appendix A Details of the datasets
A.1 Statistics and Examples
In Tables 6 and 7, we describe in numbers the four datasets under study. In Table 8, we list 16 examples from the VUA datasets, 6 examples from the LAI2009 dataset, and 4 examples from the GPT-4o-metaphors dataset.
VUA-ratings VUA-dictionary
Genre # Metaphors Lsent Novelty Score # Novel # Novel
mean | std. mean | std. >=0.5>=0.5
Fiction 3170 26.0 | 16.5 -.005 | .271 94 113
News 4712 29.9 | 14.2 .000 | .257 132 183
Academic 5499 34.9 | 16.0 .003 | .239 102 102
Conversation 1774 17.5 | 15.9 -.000 | .236 25 11
All 15155 29.4 | 16.5 .000 | .251 353 409
Table 6: Distributions and statistics of the datasets under study. # Met. is the number of metaphor words, Scorenov is the BWS novelty scores, # Nov. is the number of novel metaphors (Scorenov >= 0.5). Lsent is the length of sentences in words. We report the number of metaphors used in the sense of Nouns (N), Verbs (V) or Adjectives (Adj.) in the synthetic datasets.
Lai2009 GPT-4o-metaphors
Label # Lsent # Lsent
mean | std. mean | std.
Conventional 104 6.7 | 1.5 100 10.1 | 1.9
Novel 104 6.7 | 1.4 100 12.7 | 2.3
All 208 6.7 | 1.5 200 11.4 | 2.5
Table 7: Distributions and statistics of the datasets under study. # Met. is the number of metaphor words, Scorenov is the BWS novelty scores, # Nov. is the number of novel metaphors (Scorenov >= 0.5). Lsent is the length of sentences in words. We report the number of metaphors used in the sense of Nouns (N), Verbs (V) or Adjectives (Adj.) in the synthetic datasets.
VUA datasets
Sentence ratings dictionary direct cloze
1. ‘ Tell him I am very sorry, but I must fill the quota. ’ -0.441 conventional 9.043 7.317
2. Adam might have escaped the file memories for years, suppressed them and jerked violently <14> by those events. 0.531 novel 14.23 14.69
3. It was an excitement that <11> and I had long dreamed of that scatter of tiny, magically named islands strewn across one third of a globe. 0.278 conventional 10.17 16.08
4. The seemingly random and <11> designed to disguise a boat’s shape from the prying eyes of U-Boat captains, so it <10> in the Bahamas. 0.588 conventional 7.903 12.46
5. One Mr Clarke can not duck away from if he wants to avoid a second Winter of Discontent .. -0.094 conventional 3.165 7.454
6. This was conveniently encapsulated in the first try. 0.500 conventional 7.918 14.54
7. Thrusts of resistance ( mass demonstrations, resignations, tax rebellions, etc ) would come in crests. 0.382 novel 12.46 16.84
8. Travel: A pilgrimage sans progress Elisabeth de Stroumillo potters round Poitou 0.514 conventional 8.188 14.81
9. The Tehuana dress is by no means the most decorative variant or the closest to pre-Hispanic forms of clothing. -0.194 conventional 6.466 11.92
10. Interwoven with these images are subtler references to the metaphorical borderlines which separate Latin American <5> and North America. 0.529 conventional 10.16 12.57
11. This is often linked with a supposed denunciatory effect — the idea that the mandatory life sentence denounces murder as emphatically as possible <18> this crime. 0.294 conventional 11.02 12.61
12. He certainly held deep convictions as to the <9>, but at least a part of his apparent hostility was assumed for the occasion, a hard <7> in the end. 0.514 conventional 5.662 9.781
13. Me dad said he’s had enough Well, we were debating whether to give it to you or not. -0.633 conventional 3.752 8.038
14. Struggled with it a little 0.552 conventional 17.13 14.92
15. That’s an old trick. 0.310 conventional 4.013 11.92
16. Can you sort erm, madame out? 0.567 conventional 8.820 9.591
LAI2009
Sentence label direct cloze
17. Upon hearing the news my spirits sank conventional 4.358 11.82
18. Upon having the data my prediction sank novel 10.28 13.27
19. Those chess players are prepared for battle conventional 5.373 12.32
20. Those plastic surgeons are prepared for battle novel 6.691 13.28
21. His mental condition remains fragile conventional 7.251 13.34
22. His website popularity remains fragile novel 9.538 13.95
GPT-4o-metaphors
23. Her family was her emotional anchor during the crisis. conventional 4.322 14.12
24. The smell of coffee became an anchor to mornings that no longer came. novel 7.518 12.42
25. The software helps users navigate complex legal documents. conventional 3.830 5.575
26. He tried to navigate the silence like a sailor without stars. novel 8.041 8.402
Table 8: Examples from each dataset. The metaphor word is in boldface within sentences. For simpler presentations, we remove some words from long sentences and replace them with a tag of the number of words removed, e.g. <11>. rating is the Do Dinh et al. (2018) score, and dictionary is the Reimann and Scheffler (2024) annotation. direct is the GPT2-Base recorded surprisal, and cloze is the GPT2-Base cloze-surprisal. Paired examples are picked randomly for VUA datasets from the same genres to illustrate the differences between conventional and novel instances according to “ratings”. Sentences 1-4 are from Fiction, 5-8 from News, 9-12 from Academic and 13-16 from Conversation. Also obviously, Do Dinh et al. (2018) ratings and Reimann and Scheffler (2024) annotations do not agree in many cases.
A.2 GPT-4o-metaphors Construction
To generate the sentences in GPT-4o-metaphors, we prompt GPT-4o once for each word in the dataset. As we planned for 10 different words, we prompted the model 10 consecutive times to construct this dataset. ‘‘I am curating a dataset to be used in a study about metaphoric knowledge in pretrained language models PLMs e.g. GPT-2. The dataset should consist of sentences that correspond to target nouns/verbs. For each target noun/verb, there should be 10 sentences. The target noun/verb should be used 10 times using conventional metaphoric meanings of the target noun/verb, and 10 times using novel metaphoric meanings of the target noun/verb. Please suggest a target noun, and generate 20 sentences following the requirements above: 10 sentences with Conventional Metaphor usages, and 10 sentences with Novel Metaphor usages.’’
Appendix B Visualisations
B.1 Distributions
In Figure 3, we plot the distributions of novelty scores/labels and surprisal of the VUA datasets.
In Figure 4, we plot scatter plots between VUA-ratings scores and GPT2-base surprisal from both direct and cloze methods, grouping instances based on the location of the metaphoric word within the sentence.
Figure 3: Histograms of metaphor novelty, and direct surprisal (dark colour) of content metaphoric words (15,155) across 10 LMs. Additionally, cloze-surprisal of the same models is plotted in light colour, and surprisal from instruct-tuned variants (whenever applicable) is plotted in lined solid colour.
Figure 4: Scatter plots between metaphor novelty scores of VUA-ratings on the x-axis and surprisal (from GPT2-base) direct (blue) and cloze (red) on the y-axis at different relative positions of metaphoric words in sentences. 0-10% means metaphoric words located at the first 1/10Length1/10*Length of the sentence.